# Efficient Dynamics Modeling in Interactive Environments with Koopman Theory

### v1

@@ Abstract
The accurate modeling of dynamics in interactive environments is critical for successful long-range prediction. Such a capability could advance Reinforcement Learning (RL) and Planning algorithms, but achieving it is challenging. Inaccuracies in model estimates can compound, resulting in increased errors over long horizons. We approach this problem from the lens of Koopman theory, where the nonlinear dynamics of the environment can be linearized in a high-dimensional latent space. This allows us to efficiently parallelize the sequential problem of long-range prediction using convolution, while accounting for the agent’s action at every time step. Our approach also enables stability analysis and better control over gradients through time. Taken together, these advantages result in significant improvement over the existing approaches, both in the efficiency and the accuracy of modeling dynamics over extended horizons. We also report promising experimental results in dynamics modeling for the scenarios of both model-based planning and model-free RL.
@@ 1 Introduction
The ability to predict the outcome of an agent’s action over long horizons is a crucial unresolved challenge in Reinforcement Learning (RL)(sutton2018reinforcement, ; mnih2015human, ; Silver, ; Mnih, ). This is especially important in model-based RL and planning, where deriving a policy from the learned dynamics models allows one to efficiently accomplish a wide variety of tasks in an environment (Du2019TaskAgnosticDP, ; Hafner2020Dream, ; Silkchi2020LOOP, ; pmlr-v162-hansen22a, ; schrittwieser2020mastering, ). In fact, state-of-the-art model-free techniques also rely on dynamics models to learn a better representation for downstream value prediction tasks. Thus, obtaining accurate long-range dynamics models in the presence of input and control is crucial. In the present article, we leverage tools, techniques, and perspectives from Koopman theory koopman1931hamiltonian ; koopman1932dynamical ; brunton2021modern to address this key problem in long-range dynamics modeling in interactive environments.
@@
The application of Koopman theory allows us to linearise a nonlinear dynamical system by creating a bijective mapping to linear dynamics in a possibly infinite dimensional space of observables. [^1] The term observables can be misleading as it refers to the latent space in the machine learning jargon. Conveniently, a deep neural network can learn to produce such a mapping, enabling a reliable approximation of the non-linear dynamics in the finite dimension of a linear latent space so that only a finite subset of the most relevant (complex-valued) Koopman observables need to be tracked.
We show that this linearization has two major benefits: 1) The eigenvalues of the linear latent-space operator are directly related to the stability and expressive power of the controlled dynamics model. Furthermore, spectral Koopman theory provides a unifying perspective over several existing directions, including long-range sequence modeling using state-space models gu2022on , and the use of unitary matrices arjovsky2016unitary in latent dynamic models; 2) Linearization can help to avoid a computational bottleneck due to the sequential nature of dynamics. This is achieved by diagonalization of the linear operator and a reformulation using convolution. In other words, one can perform efficient parallel training of the model over time steps, despite dealing with a controlled dynamical system.
Our experimental results in offline-RL datasets demonstrate the effectiveness of our approach for reward and state prediction over a long horizon. In particular, we report competitive results against sequentially-trained models that use MLP for every time-step, as well as in comparison to transformer models. We also present encouraging results for model-based planning and model-free RL, where our use of Koopman theory for dynamics modeling improves representation learning.
@@ 2.1 Koopman Theory for Dynamical Systems
The infinite dimensionality of the Koopman operator is a limitation in practice, and thus an invariant subspace $\mathcal{G}\subset\mathcal{F}$ is often used, so that $\mathcal{K}g\in\mathcal{G}$. $\mathcal{G}$ is spanned by a finite set of observables ${g_{1},\ldots,g_{m}}$, where often we assume $m\gg n$.
Constraining the Koopman operator on this invariant subspace results in a finite-dimensional linear operator $K\in\mathds{C}^{m\times m}$, called the Koopman matrix, which satisfies $g(x_{t+1})=Kg(x_{t}).$ Usually, the base observation functions $\{g_{i}\}_{i}$ are hand-crafted using knowledge of the system’s underlying physics. However, data-driven methods have recently been proposed to learn the Koopman operator by representing the base observations or their eigenfunction basis using deep neural networks, where a decoder reconstructs the input from linear latent dynamics (e.g.lusch2018deep, ; champion2019data, ). Linearizing the dynamics allows for closed-form solutions for the predictions of the dynamical system, as well as stability analysis based on the eigendecomposition of the Koopman matrix.
@@
By using a diagonal Koopman matrix, we are effectively offloading the task of learning the suitable eigenfunctions of the Koopman operator that form an invariant subspace to the neural network encoder. Additionally, for a continuous time input, to be able to model higher-order frequencies in the eigenspectrum and provide better approximations, one could adaptively choose the most relevant eigenvalues and eigenfunctions for the current state Lusch2017DeepLF . This means $g(x_{t+1})=K_{D}(\lambda(g(x_{t})))g(x_{t}),$ where $\lambda:\mathcal{G}\to\mathds{C}^{m}$ can be a neural network.
@@ 2.2 Approximate Koopman with Control Input
The Koopman operator can be extended to non-linear control systems, where the state of the system is influenced by an external control input $u_{t}$ such that $x_{t+1}=F(x_{t},u_{t})$ or $\frac{dx}{dt}=f(x,u)$. Simply treating the pair $(x,u)$ as the input $x$, the Koopman operator becomes $(\mathcal{K}g)(x_{t},u_{t})=g(F(x_{t},u_{t}),u_{t+1})=g(x_{t+1},u_{t+1}).$ For a control affine setting, where
$f(x,u)=f_{0}(x)+\sum_{i=1}^{m}f_{i}(x)u_{i}$
one could show that the Koopman operator is bilinearized brunton2021modern :
@@
More generally, one could make the Koopman operator a function of $u_{t}$, so that the resulting Koopman matrix satisfies $g(x_{t+1})=K(u_{t})g(x_{t})$. This is the approach taken by pmlr-v162-weissenbacher22a to model symmetries of dynamics in offline RL. When combined with the diagonal Koopman matrix of Eq. 1, and fixed modulus $|\lambda_{i}|=1\,\forall i$, the Koopman matrix becomes a product of $SO(2)$ representations, and the resulting latent dynamics model, resembles a special setting of the symmetry-based approach of pmlr-v162-mondal22a , where group representations are used to encode states and state-dependent actions.
An alternate approach to simplifying the Koopman operator with a control signal assumes a decoupling of state and control observables $g(x,u)=[g(x),f(u)]=[g_{1}(x),...,g_{m}(x),f_{1}(u),...,f_{n}(u)]$. This gives rise to a simple linear evolution:
@@
where $K\in\mathds{C}^{m\times m}$ and $L\in\mathds{C}^{m\times l}$ are matrices representing the linear dynamics of the state observables and control input, respectively. We build on this approach in our proposed methodology. Although, as shown in brunton2016discovering ; bruder2019modeling , even assuming a linear control term $Lu_{t}$ can perform well in practice, we use neural networks for both $f$ and $g$. Our rationale for choosing this formulation is that the additive form of Eq. 2 combined with the diagonalized Koopman matrix of Eq. 1, enable fast parallel training across time-steps using a convolution operation where the convolution kernel can be computed efficiently. Moreover, this setting is amenable to the analysis of gradient behaviour, as discussed in Section 3.2.1.
2.3 Reinforcement Learning
Value-based RL algorithms have two steps of policy evaluation and policy improvement. For example, $Q$-learning estimates the optimal action-value function, denoted by $Q^{*}(s,a)$, that represents the maximum expected cumulative reward achievable by taking action $a$ in state $s$ and improves the policy by setting it to $arg\max_{a}Q^{*}(s,a)$ (mnih2015human, ; MnihKSGAWR13, ). A policy-based RL algorithm directly optimizes the policy to maximize the expected return. Policy gradient methods update the policy by adjusting the policy function’s parameters by estimating the gradient of $J$. Actor-critic methods combine Q-learning and policy gradient ones, such that the Q-network (critic) learns the action value under the current policy, and the policy network (actor) tries to maximize it (haarnoja2018soft, ; LillicrapHPHETS15, ; Schulman2017PPO, ).
2.4 Forward dynamics modeling in RL
Dynamics models have been instrumental in attaining impressive results on various tasks such as Atari (schrittwieser2020mastering, ; Hafner2020Dream, ) and continuous control (Hafner2020Dream, ; Janner2019MBPO, ; sikchi2021learning, ; Lowrey2018POLO, ). By building a model of environment dynamics, model-based RL can generate trajectories used for training the RL algorithm. This can significantly reduce the sample complexity in comparison to model-free techniques. However, in model-based RL, inaccurate long-term predictions can generate poor trajectory rollouts and lead to incorrect expected return calculations, resulting in misleading policy updates. Forward dynamics modeling has also been successfully applied to model-free RL to improve the sample efficiency of the existing model-free algorithms. These methods use it to design self-supervised auxiliary losses for representation learning using consistency in forward dynamics in the latent and the observation space jaderberg2017reinforcement ; Schwarzer2020DataEfficientRL ; Srinivas2020CURLCU .
2.5 Model-based Planning
where $\gamma$ is typically set to 1, that is, there is no discounting. Heuristic population-based methods, such as a cross entropy method rubinstein2004cross are often used for dynamic planning. These methods perform a local trajectory optimization problem, corresponding to the optimization of a cost function up to a certain number of time steps in the future (Williams2016MPPI, ; Williams2017InfoMPPI, ; Okada2019VarMPC, ). In contrast to Q-learning, this is myopic and can only plan up to a certain horizon, which is predetermined by the algorithm.
One can also combine MPC with RL to approximate long-term returns of trajectories that can be calculated by bootstrapping the value function of the terminal state. In particular, methods like TD-MPC pmlr-v162-hansen22a and LOOP Silkchi2020LOOP combine value learning with planning using MPC. The learned value function is used to bootstrap the trajectories that are used for planning using MPC. Additionally, they learn a policy using the Deep Deterministic Policy Gradient Lillicrap2015ContinuousCW or Soft Actor-Critic (SAC) haarnoja2018soft to augment the planning algorithm with good proposal trajectories. Alternative search method such as Monte Carlo Tree search (coulom2006efficient, ) is used for planning in discrete action spaces. The performance of the model-based planning method heavily relies on the accuracy of the learned model.
@@ 3.1 Linear Latent Dynamics Model
Our task in dynamics modeling is to predict the sequence of future states ${s_{t+1},....,s_{t+\tau}}$ given a starting state $s_{t}$ and some action sequence ${a_{t},a_{t+1},...,a_{t+\tau-1}}$, assuming a Markovian system. Following the second approach described in Section 2.2, we assume that state and control observables are decoupled, and use neural network encoders to encode both states and actions
@@
and the solution is given by $x(s)=e^{Ks}x(0)+\int_{0}^{s}e^{K(s-t)}Lu(t)\mathrm{d}t$ where $e^{Ks}$ is a matrix exponential. One could discretize this to get $\hat{x}_{t+1}=\bar{K}x_{t}+\bar{L}u_{t}$, where $\bar{K}$ and $\bar{L}$ are obtained by Zero-Order Hold (ZOH) discretization of the continuous-time equation iserles2009first :
@@ 3.2 Diagonalization, Efficiency and Stability of the Koopman Operator
To unroll the latent space using Eq. 10 we need to perform matrix exponentiation and dense multiplication. Using a diagonal Koopman matrix $\bar{K}=\mathrm{diag}(\bar{\lambda}_{1},\ldots,\bar{\lambda}_{m})$, this calculation is reduced to computing an $m\times(\tau+1)$ complex-valued Vandermonde matrix:
$\Lambda=\begin{bmatrix}1&\bar{\lambda}_{1}&\bar{\lambda}_{1}^{2}&\dots&\bar{\lambda}_{1}^{\tau}\\ 1&\bar{\lambda}_{2}&\bar{\lambda}_{2}^{2}&\dots&\bar{\lambda}_{2}^{\tau}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&\bar{\lambda}_{m}&\bar{\lambda}_{m}^{2}&\dots&\bar{\lambda}_{m}^{\tau}\end{bmatrix}$ (12)
and doing a row-wise circular convolution of this matrix with a sequence of vectors. Recall that $x,c\in\mathds{C}^{m}$ are complex vectors. We use superscript to index their elements – i.e.$x_{t}=[x_{t}^{1},\ldots,x_{t}^{m}]$. With this notation, the $i$-th element of the predicted Koopman observables for future time steps is given by (see Appendix B for the derivation):
@@ 3.2.1 Gradients Through Time
We show how we can control the behavior of gradients through time by constraining the real part of the eigenvalues of the Diagonal Koopman matrix. Let $\mu$ and $\omega$ refer to the real and imaginary part of the Koopman eigenvalues – that is $\bar{\lambda}_{j}=\mu_{j}+i\omega_{j}$.
@@ 3.2.2 Initialization of the Eigenspectrum
From Theorem A.1, we know that the real part of the eigenvalues impacts the gradient’s behavior through time. To avoid vanishing gradients and account for prediction errors over longer horizons, one could eliminate the real part $\mu_{j}:=0$. This choice turns the latent transformations into blocks of 2D rotations mondal2022eqr . This is also related to using Unitary Matrices to avoid vanishing or exploding gradients in Recurrent Neural Networks arjovsky2016unitary . However, intuitively, we might prefer to prioritize closer time steps. This can be done using small negative values e.g.$\mu_{j}\in\{-0.1,-0.2,-0.3\}$. An alternative to having a fixed real part is to turn it into a bounded learnable parameter $\mu_{j}\in[-0.3,-0.1]$. We empirically found $\mu_{j}:=-0.2\;\forall j$, and $\omega_{j}:=j\pi$ to be good choices and use this as our default initialization. In Appendix D we report the results of an ablation study with different initialization schemes.

### v4

@@ Abstract
The accurate modeling of dynamics in interactive environments is critical for successful long-range prediction. Such a capability could advance Reinforcement Learning (RL) and Planning algorithms, but achieving it is challenging. Inaccuracies in model estimates can compound, resulting in increased errors over long horizons. We approach this problem from the lens of Koopman theory, where the nonlinear dynamics of the environment can be linearized in a high-dimensional latent space. This allows us to efficiently parallelize the sequential problem of long-range prediction using convolution while accounting for the agent’s action at every time step. Our approach also enables stability analysis and better control over gradients through time. Taken together, these advantages result in significant improvement over the existing approaches, both in the efficiency and the accuracy of modeling dynamics over extended horizons. We also show that this model can be easily incorporated into dynamics modeling for model-based planning and model-free RL and report promising experimental results.
@@ 1 Introduction
The ability to predict the outcome of an agent’s action over long horizons is a crucial unresolved challenge in Reinforcement Learning (RL) (Sutton & Barto, 2018, Mnih et al., 2015, Silver et al., 2017, Mnih et al., 2016). This is especially important in model-based RL and planning, where deriving a policy from the learned dynamics models allows one to efficiently accomplish a wide variety of tasks in an environment (Du & Narasimhan, 2019, Hafner et al., 2020, Sikchi et al., 2021, Hansen et al., 2022, Schrittwieser et al., 2020, Jain et al., 2022). In fact, state-of-the-art model-free techniques also rely on dynamics models to learn a better representation for downstream value prediction tasks (Schwarzer et al., 2020). Thus, obtaining accurate long-range dynamics models in the presence of input and control is crucial. In this work, we leverage techniques and perspectives from Koopman theory (Koopman, 1931, Mauroy et al., 2020, Koopman & Neumann, 1932, Brunton et al., 2021) to address this key problem in long-range dynamics modeling of interactive environments. The application of Koopman theory allows us to linearise a nonlinear dynamical system by creating a bijective mapping to linear dynamics in a possibly infinite dimensional space of observables. [^1] The term observables can be misleading as it refers to the latent space in the machine learning jargon.
@@
Conveniently, a deep neural network can learn to produce such a mapping, enabling a reliable approximation of the non-linear dynamics in the finite dimension of a linear latent space so that only a finite subset of the most relevant (complex-valued) Koopman observables need to be tracked.
We show that this linearization has two major benefits: 1) The eigenvalues of the linear latent-space operator are directly related to the stability and expressive power of the controlled dynamics model. Furthermore, spectral Koopman theory provides a unifying perspective over several existing directions on stabilizing gradients through time, including long-range sequence modeling using state-space models (Gu et al., 2020; 2022b; 2022a, Gupta et al., 2022), and the use of unitary matrices (Arjovsky et al., 2016) in latent dynamic models; 2) This can help to avoid a computational bottleneck due to the sequential nature of dynamics. This is achieved by a reformulation using convolution and diagonalization of the linear operator. In other words, one can perform efficient parallel training of the model over time steps, despite dealing with a controlled dynamical system.
Our experimental results in offline-RL datasets demonstrate the effectiveness of our approach for reward and state prediction over a long horizon. In particular, we report competitive results against dynamics modeling baselines that use Multi-Layer Perceptrons (MLPs), Gated Recurrent Units (GRUs), Transformers and Diagonal State Space Models Gu et al. (2022a) while being significantly faster. Finally, we also present encouraging results for model-based planning and model-free RL with our Koopman-based dynamics model.
In summary, our contributions are:
  1. 1.
    We formulate an efficient linear latent dynamics model from Koopman theory that can be parallelized during training for long-range dynamics modeling in interactive environments.
  2. 2.
    We demonstrate the use of the spectral decomposition of the Koopman matrix, which enables better control over gradients through time and accelerated training.
  3. 3.
    We show that our technique serves as a preferred alternative for existing dynamics models in planning and RL algorithms, offering enhanced performance, speed and training stability.
@@ 2.1 Koopman Theory for Dynamical Systems
The infinite dimensionality of the Koopman operator presents practical limitations. Addressing this, we seek to identify a subspace $\mathcal{G}\subset\mathcal{F}$, spanned by a finite set of observables ${g_{1},\ldots,g_{m}}$ where typically $m\gg n$, that approximates invariance under the Koopman operator.
Constraining the Koopman operator on this invariant subspace results in a finite-dimensional linear operator $K\in\mathds{C}^{m\times m}$, called the Koopman matrix, which satisfies $g(x_{t+1})=Kg(x_{t}).$ Usually, the base observation functions $\{g_{i}\}_{i}$ are hand-crafted using knowledge of the system’s underlying physics. However, data-driven methods have recently been proposed to learn the Koopman operator by representing the base observations or their eigenfunction basis using deep neural networks, where a decoder reconstructs the input from linear latent dynamics (e.g.,  Lusch et al., 2018, Champion et al., 2019). Linearizing the dynamics allows for closed-form solutions for the predictions of the dynamical system, as well as stability analysis based on the eigendecomposition of the Koopman matrix Fan et al. (2022), Yi & Manchester (2023).
@@
By using a diagonal Koopman matrix, we are effectively offloading the task of learning the suitable eigenfunctions of the Koopman operator that form an invariant subspace to the neural network encoder. Additionally, for a continuous time input, to be able to model higher-order frequencies in the eigenspectrum and provide better approximations, one could adaptively choose the most relevant eigenvalues and eigenfunctions for the current state (Lusch et al., 2018). This means $g(x_{t+1})=K_{D}(\lambda(g(x_{t})))g(x_{t}),$ where $\lambda:\mathcal{G}\to\mathds{C}^{m}$ can be a neural network.
@@ 2.2 Approximate Koopman with Control Input
The Koopman operator can be extended to non-linear control systems, where the state of the system is influenced by an external control input $u_{t}$ such that $x_{t+1}=F(x_{t},u_{t})$ or $\frac{dx}{dt}=f(x,u)$. Simply treating the pair $(x,u)$ as the input $x$, the Koopman operator becomes $(\mathcal{K}g)(x_{t},u_{t})=g(F(x_{t},u_{t}),u_{t+1})=g(x_{t+1},u_{t+1}).$ If the effect of the control input on the system’s dynamics is linear, i.e., the control affine setting, we have
$f(x,u)=f_{0}(x)+\sum_{i=1}^{m}f_{i}(x)u_{i}.$
assuming a finite number of eigenfunctions. For such control-affine system, one could show that the Koopman operator is bilinearized (Brunton et al., 2021, Bruder et al., 2021):
@@
More generally, one could make the Koopman operator a function of $u_{t}$, so that the resulting Koopman matrix satisfies $g(x_{t+1})=K(u_{t})g(x_{t})$. This is the approach taken by Weissenbacher et al. (2022) to model symmetries of dynamics in offline RL. When combined with the diagonal Koopman matrix of Eq. 1, and fixed modulus $|\lambda_{i}|=1\,\forall i$, the Koopman matrix becomes a product of $SO(2)$ representations, and the resulting latent dynamics model, resembles a special setting of the symmetry-based approach of Mondal et al. (2022), where group representations are used to encode states and state-dependent actions.
An alternate approach to approximating the Koopman operator with a control signal assumes a decoupling of state and control observables $g(x,u)=[g(x),f(u)]=[g_{1}(x),...,g_{m}(x),f_{1}(u),...,f_{n}(u)]$ Brunton et al. (2021), Bruder et al. (2019). This gives rise to a simple linear evolution:
@@
where $K\in\mathds{C}^{m\times m}$ and $L\in\mathds{C}^{m\times l}$ are matrices representing the linear dynamics of the state observables and control input, respectively. We build on this approach in our proposed methodology. Although, as shown in Brunton et al. (2016; 2021), Bruder et al. (2019), even assuming a linear control term $Lu_{t}$ can perform well in practice, we use neural networks for both $f$ and $g$. Our rationale for choosing this formulation is that the additive form of Eq. 2 combined with the diagonalized Koopman matrix of Eq. 1, enable fast parallel training across time-steps using a convolution operation where the convolution kernel can be computed efficiently. Moreover, this setting is amenable to the analysis of gradient behaviour, as discussed in Section 3.2.1. For a more comprehensive overview of Koopman theory in the context of controlled and uncontrolled dynamics, we direct readers to Brunton et al. (2021), Mauroy et al. (2020).
@@ 3 Proposed Model
Figure 2: A schematic of the latent Koopman dynamics model. Both actions and initial state embedding are encoded into a latent space in complex ($\mathbb{C}$) domain before passing through the Koopman dynamics block. (see Appendix J for an efficient Jax implementation of the model)
@@ 3.1 Linear Latent Dynamics Model
Our task in dynamics modeling is to predict the sequence of future states ${s_{t+1},....,s_{t+\tau}}$ given a starting state $s_{t}$ and some action sequence ${a_{t},a_{t+1},...,a_{t+\tau-1}}$, assuming a Markovian system. Following the second approach described in Section 2.2, we assume that state and control observables are decoupled, and use neural network encoders [^2] We use a CNN for pixel input and an MLP for state-based input and actions. to encode both states and actions
@@
and the solution is given by $x(s)=e^{Ks}x(0)+\int_{0}^{s}e^{K(s-t)}Lu(t)\mathrm{d}t$ where $e^{Ks}$ is a matrix exponential. One could discretize this to get $\hat{x}_{t+1}=\bar{K}x_{t}+\bar{L}u_{t}$, where $\bar{K}$ and $\bar{L}$ are obtained by Zero-Order Hold (ZOH) discretization of the continuous-time equation (Iserles, 2009):
@@ 3.2 Diagonalization, Efficiency and Stability of the Koopman Operator
To unroll the latent space using Eq. 5 we need to perform matrix exponentiation and dense multiplication. Using a diagonal Koopman matrix $\bar{K}=\mathrm{diag}(\bar{\lambda}_{1},\ldots,\bar{\lambda}_{m})$, this calculation is reduced to computing an $m\times(\tau+1)$ complex-valued Vandermonde matrix $\Lambda$, where $\Lambda_{i,j}=\bar{\lambda}_{i}^{j}$, and doing a row-wise circular convolution of this matrix with a sequence of vectors.
Recall that $x,c\in\mathds{C}^{m}$ are complex vectors. We use superscript to index their elements – i.e.$x_{t}=[x_{t}^{1},\ldots,x_{t}^{m}]$. With this notation, the $i$-th element of the predicted Koopman observables for future time steps is given by (see Appendix B for the derivation):
@@ 3.2.1 Gradients Through Time
We show how we can control the behavior of gradients through time by constraining the real part of the eigenvalues of the Diagonal Koopman matrix. Let $\mu$ and $\omega$ refer to the real and imaginary part of the Koopman eigenvalues – that is $\lambda_{j}=\mu_{j}+i\omega_{j}$.
@@ 3.2.2 Initialization of the Eigenspectrum
From Theorem A.1, we know that the real part of the eigenvalues impacts the gradient’s behavior through time. To avoid vanishing gradients and account for prediction errors over longer horizons, one could eliminate the real part $\mu_{j}:=0$. This choice turns the latent transformations into blocks of 2D rotations (Mondal et al., 2022). This is also related to using Unitary Matrices to avoid vanishing or exploding gradients in Recurrent Neural Networks (Arjovsky et al., 2016). However, intuitively, we might prefer to prioritize closer time steps. This can be done using small negative values e.g.$\mu_{j}\in\{-0.1,-0.2,-0.3\}$. An alternative to having a fixed real part is to turn it into a bounded learnable parameter $\mu_{j}\in[-0.3,-0.1]$. We empirically found $\mu_{j}:=-0.2\;\forall j$, and $\omega_{j}:=j\pi$ to be good choices and use this as our default initialization. In Appendix H we report the results of an ablation study with different initialization schemes.